itskaelen:

A metal washer is made by cutting out a circle 13.2 mm in diameter from a larger circle. To the nearest tenth, what is the area of the washer?

loveboy:

Diameter=13.2 mm Radius = 13.2/2=6.6 mm Area= pi* (radius)^2 Area= pi*(6.6)^2=136.84 so according to nearest tenth, area=136.8 mm^2